A 50-gallon electric water heater is used to heat tap water with an initial temperature of 15oC. The heating level is set at a maximum while several people are taking their showers. If the maximum heating level of the heater is 7.5kW of electricity and the water use rate is 3 gal/min, what is the temperature of the water exiting the heater? Assume that the system is at steady state and the heater is 100% efficient; that is it is perfectly insulated and all of the energy used goes to heat the water. If 10% of the heat is lost to the surrounding what is the temperature of the exiting water?
Answer #1

mass m = 3 gallons x P = (3x 3.7811) xlkg/ = 0-189 kals min 60 second heat input q = 7.5 kW 1000J/s) =7500J/s Ikw q=m COT whe

sir,please give a rating